# Improving Deep Reinforcement Learning by Reducing the Chain Effect of Value and Policy Churn

### v1

@@ 6.2 Results for CHAIN PPO in MuJoCo and DMC
Corresponding to the second DRL scenario discussed in Section 5.1, we focus on the policy churn in Proximal Policy Optimization (PPO) (Schulman et al., 2017), and try to answer the first three questions for policy gradient-based RL algorithms. We build the experiments on the public implementation of PPO from CleanRL (Huang et al., 2022) and use the continuous control tasks in MuJoCo and DeepMind Control (DMC) as the environments for evaluation. Following the same principle, we implement CHAIN PPO by adding the policy churn reduction regularization to the standard PPO policy training (Eq. 6), with no other modification to the public PPO implementation.
First, to understand the level of churn while PPO is training, we compare PPO and CHAIN PPO with different choices of $\lambda_{\pi}$ in terms of policy churn and episode return. In summary, we observed that PPO also exhibits clear policy churn and CHAIN significantly reduces it throughout learning, which answers Question (1). Figure 11 shows the details for this analysis. Note that more policy churn makes it more likely to violate the trust region as $\mathcal{C}_{\pi}\propto r(\phi,\phi^{\prime})$ discussed in Section 5.1. In turn, we also observed CHAIN PPO consistently outperforms PPO in Ant-v4 and HalfCheetah-v4 across different choices of $\lambda_{\pi}$.
Figure 5: The evaluation of CHAIN PPO in MuJoCo regarding episode return. Curves and shades denote means and standard errors over twelve random seeds.
Further, we aim to answer Question (2) and evaluate whether CHAIN can improve the learning performance of PPO in terms of episode return. For CHAIN PPO (Auto), we set the target relative loss scale $\beta$ to 0.1 for MuJoCo tasks and 0.02 for DMC tasks. The results are reported in Figure 5 and Figure 6 for MuJoCo and DMC, respectively. The results show that CHAIN PPO outperforms PPO in most cases with higher sample efficiency and final episode return, often significantly. We believe that our results reveal a promising direction to improve more trust-region-based and constraint-based methods in DRL by addressing the issues caused by churn.
Figure 6: The evaluation of CHAIN PPO in DeepMind Control (DMC) tasks regarding episode return. Curves and shades denote means and standard errors over twelve random seeds.
@@ 6.3 Results for Deep Actor-Critic Methods with CHAIN in MuJoCo and D4RL
Online AC methods
We use TD3 (Fujimoto et al., 2018) and SAC (Haarnoja et al., 2018) and MuJoCo environments based on the public implementation of TD3 and SAC from CleanRL. Since AC methods are affected by churn in a bilateral manner, we consider two variants of CHAIN, either of which only applies the value churn reduction (VCR) or the policy churn reduction (PCR).
For Question (1), we again find that both TD3 and SAC exhibit value and policy churn in all environments, and CHAIN-VCR and CHAIN-PCR effectively reduce them respectively in Figure 12 and 13. For Question (2), Figure 7 shows the evaluation results regarding episode return. We can see that CHAIN-PCR often improves the learning performance, especially for Ant-v4; in contrast, CHAIN-VCR improves slightly. We hypothesize that this is because the policy interacts with the environment directly, and the target critic-network also helps to reduce the value churn due to its delayed synchronization with the online critic.
Due to the limitation of space, we refer the readers to Appendix C.3 for more results on churn reduction, the influence of different choices of $\lambda_{Q},\lambda_{\pi}$, the results of combining VCR and PCR, as well as the effect of auto-adjustment of the regularization coefficient for TD3 and SAC.
@@
Offline AC methods
In Offline RL, a policy is trained over a fixed dataset. We investigate if reducing churn can also improve the convergence of Offline RL. We use IQL (Kostrikov et al., 2022) with D4RL Antmaze dataset (Fu et al., 2020) and AWAC (Nair et al., 2020) with Adroit for our demonstration due to their popularity and good performance in corresponding tasks. Concretely, we use the public implementation and benchmark scores for IQL and AWAC from CORL [^1] https://github.com/tinkoff-ai/CORL. To apply CHAIN to IQL and AWAC, we implement the regularization for value churn reduction (VCR, Eq. 7) and policy churn reduction (PCR, Eq. 8) separately by adding a couple of lines of code without any other modification. We use $\lambda_{\pi}=1e3$ for both CHAIN IQL (PCR) and CHAIN AWAC (PCR); $\lambda_{Q}=0.01$ for CHAIN IQL (VCR) and $0.1$ for CHAIN AWAC (VCR) across different tasks. The results are summarized in Table 1 and 2.
@@
We observe that both CHAIN PCR and CHAIN VCR improve the scores for IQL and AWAC in most Antmaze and Adroit tasks. We hypothesize that CHAIN suppresses churn in the training of value and policy networks, thus reducing the bias caused by churn in parameter updates. One thing here that differs from TD3 and SAC considered in the online setting is that, the policy network of IQL has no impact on the training of the value networks, since the value networks (i.e., $Q$ and $V$) are trained purely based on in-sample data without accessing $a^{\prime}=\pi_{\phi}(s^{\prime})$. Thus, although IQL does not exhibit a chain effect explicitly, the policy and value networks of IQL still have churns, which are reduced by CHAIN in this case. We provide a further empirical study in Appendix C.4.
@@ 6.4 Scaling DRL Agents with CHAIN
It is widely known that scaling DRL agents up is challenging. Naively scaling DRL agents by widening or deepening the conventional MLP networks straightforwardly often fails and could even lead to collapse. Here, we investigate the relationship between churn and scale for DRL agents, as well as the effect of CHAIN on boosting scaling performance, in order to answer Question (3).
We take PPO and MuJoCo tasks as the exemplary setting and scale up both the policy and value networks by a scale-up ratio within $\{2,4,8,16\}$ via widening or deepening. Note that the default network architecture (i.e., when the scale-up ratio equals one) for both the policy and value networks is a two-layer MLP with 256 neurons for each layer, followed by an output layer. As expected, we observed that the performance of PPO degraded severely as the scale-up ratio increased. This is shown by the solid gray lines in Figure 8. Inspired by the prior study (Obando-Ceron et al., 2024), we found using a decreased learning rate as lr / sqrt(scale-up ratio) alleviates the degradation of PPO scaling to some degree (shown by the solid red lines). We then use the learning rate setting below by default. From the perspective of churn, we also observed that scaling PPO escalates the scale of the churn in PPO’s policy. More results can be found in Appendix C.5.
(a) PPO Scaling by Widening
(b) PPO Scaling by Deepening
Driven by the relationship between churn and scaling observed in our experiments, we then evaluate the effect of CHAIN in this scaling setting. The results are shown in Figure 8 with dashed lines. By comparing the lines in the same color, we found that CHAIN improves the learning performance of PPO across almost all scale-up ratios and the two learning rate settings.
Table 3: Scaling PPO with CHAIN. The means and standard errors of final episode returns over six seeds are reported.
@@
In addition, we extend the training horizon from 2M to 10M for Ant-v4, Humanoid-v4, and Walker2d-v4. The results are reported in Table 3. For both widening or deepening cases, CHAIN helps to scale up PPO and achieves clear improvement in terms of episode return. Comparatively, scaling by widening slightly outperforms that by deepening, which aligns with the observation in prior work (Ota et al., 2020) to some extent.
Our results indicate that uncontrolled churn could be a reason for the notorious scaling issue of DRL agents, and CHAIN helps the scaling by reducing churn effectively. Though somewhat appealing, CHAIN is insufficient to fully address the scaling issue of DRL agents as degradation still exists when using a large scale-up ratio.
7 Conclusion and Discussion
In this paper, we conduct a formal study of churn in a general view and present the chain effect of value and policy churn. The chain effect indicates a degenerating cycle, which biases parameter updates throughout learning. We propose an easy-to-implement method for value and policy churn reduction. Our experimental results demonstrate the effectiveness of our method in reducing churn and improving learning performance in a range of DRL scenarios.
However, our work is limited in the directions below. First, the theoretical analysis of the chain effect under concrete assumptions remains to be explored. Other perspectives that may influence churn, such as network structure, representation learning, and experience replay are not considered in this work, which are worthwhile to study in the future. Moreover, although we provided a simple method to adjust the regularization coefficients dynamically throughout learning by keeping a consistent relative loss scale, it is not sufficient for us to use the same hyperparameter for different domains. We believe that using normalization techniques to unify the scales in different domains is necessary to address this point, similar to the work in (Hafner et al., 2023). Besides, for algorithms that involve the learning of both policy and value (e.g., deep AC methods), the implementation of CHAIN in different problems faces the question of choosing the best option among PCR, VCR and DCR. For this, we expect to develop a better method to integrate the effect of PCR and VCR in the future. Another remaining problem is the lack of an in-depth understanding of the positive and negative effects of churn on the generalization of DRL agents, which could drive new methods that better leverage the potential of churn.
Appendix A Additional Formal Analysis
A.1 The NTK Expressions of Two Types of Deviation
@@
The NTK expression of the two types of deviation above indicates that action gradient deviation and policy value deviation are mainly influenced by the second-order partial derivatives of $Q_{\theta}$ and the deterministic policy gradient (DPG) of $\pi_{\phi}$ at the state $\bar{s}$ considered, in addition to the parameter update. An interesting observation here is, policy value deviation of policy churn has an implicit optimization effect (while altered by $\Delta_{\phi}$) for the reference states (i.e., $\bar{s}$) although the parameter update $\phi\rightarrow\phi^{\prime}$ is performed for the training states (i.e., $s$). We observe that $\mathcal{D}^{\pi}_{Q}$ is positive in overall in our empirical investigation later (Appendix C). However, careful considerations are needed in the future because the implicit optimization could be delusional as churn also occurs in $Q_{\theta}$ through the learning process.
A.2 Complete Discussions and Derivations for Section 4.3
@@
Appendix B Experimental Details
B.1 Compute Resources and Time Cost
@@
B.2 Empirical Metrics for the Investigation of the Chain Effect
@@
We use $N=50$ and $N=20$ for MuJoCo and MinAtar environments and compute the metrics at an interval of 1k parameter updates. For each type of metric in Table 4 and update number $i\in\{1,\dots,N\}$, we compute the mean of the quantities throughout learning. We refer the readers to Figure 10, 12 and 13 for the results in Appendix C.
B.3 Code Implementation
@@ B.4 Other Discussions
Churn Reduction Hyperparameters (refer to the study in Fig. 7, 12, 13)
@@
Appendix C Complete Results
C.1 More Empirical Analysis on the Value Churn in DoubleDQN
@@
C.2 More Results of CHAIN PPO
@@
C.3 More Results of CHAIN for Deep AC Methods
@@
Moreover, we observe a positive policy value deviation of action churn ($\hat{\mathcal{D}}^{\pi}_{Q}$), which matches our discussion on implicit optimization effect of the policy churn in Appendix A; in contrast, the value churn ($\hat{\mathcal{C}}_{Q_{sa}}$) fluctuates above and below 0.
@@
Figure 17 shows the evaluation of the automatic adjustment method of $\lambda_{\pi}$ for CHAIN TD3 in four MuJoCo tasks. We can observe that CHAIN (Auto) achieves comparable performance with manually selected coefficients and improves TD3, similarly to the conclusions we found in Figure 45 and 6.
@@
C.4 More Results of CHAIN for IQL with Sequential Training
@@
We report the final scores of IQL (sequential) with means and standard errors over 12 seeds in Table 8. The results show that IQL (sequential) performs worse than IQL in 5 of 6 tasks. The difference between IQL (sequential) and IQL can be fully attributed to the difference in the training dynamics, mainly on the policy network. To provide some possible explanation, we guess that the difference in the training dynamics of the policy network of IQL results in a further difference in the policy output on out-of-sample states. However, it is somewhat tricky to explain the difference between IQL (sequential) and IQL. The dynamics is beyond the scope of the chain effect of churn mechanism studied in our work.
@@
C.5 More Results for Scaling PPO with CHAIN

### v2

@@ 6.2 Results for CHAIN PPO in MuJoCo and DMC
Corresponding to the second DRL scenario discussed in Section 5.1, we focus on the policy churn in Proximal Policy Optimization (PPO) (Schulman et al., 2017) and try to answer the first three questions for policy gradient-based RL algorithms. We build the experiments on the public implementation of PPO from CleanRL (Huang et al., 2022) and use the continuous control tasks in MuJoCo and DeepMind Control (DMC) as the environments for evaluation. Following the same principle, we implement CHAIN PPO by adding the policy churn reduction regularization to the standard PPO policy training (Eq. 6), with no other modification to the public PPO implementation.
First, to understand the level of churn, we compare PPO and CHAIN PPO with different choices of $\lambda_{\pi}$ in terms of policy churn and episode return. In summary, we observed that PPO also exhibits clear policy churn, and CHAIN significantly reduces it throughout learning, which answers Question (1). Figure 10 shows the details for this analysis. Note that more policy churn makes it more likely to violate the trust region as $\mathcal{C}_{\pi}\propto r(\phi,\phi^{\prime})$ discussed in Section 5.1. In turn, we also observed CHAIN PPO consistently outperforms PPO in Ant-v4 and HalfCheetah-v4 across different choices of $\lambda_{\pi}$.
Further, we aim to answer Question (2) and evaluate whether CHAIN can improve the learning performance of PPO in terms of episode return. For CHAIN PPO (Auto), we set the target relative loss scale $\beta$ to 0.1 for MuJoCo tasks and 0.02 for DMC tasks. The results for MuJoCo and DMC tasks are reported in Figure 5. The results show that CHAIN PPO outperforms PPO in most cases with higher sample efficiency and final episode return, often significantly. We believe that our results reveal a promising direction to improve more trust-region-based and constraint-based methods in DRL by addressing the issues caused by churn.
Figure 5: The evaluation of CHAIN PPO in MuJoCo and DeepMind Control (DMC) tasks regarding episode return. Curves and shades denote means and standard errors over twelve random seeds.
@@ 6.3 Results for Deep Actor-Critic Methods with CHAIN in MuJoCo and D4RL
Online AC methods We use TD3 (Fujimoto et al., 2018) and SAC (Haarnoja et al., 2018) and MuJoCo environments based on the public implementation of TD3 and SAC from CleanRL. Since AC methods are bilaterally affected by churn, we consider two variants of CHAIN, either of which only applies the value churn reduction (VCR) or the policy churn reduction (PCR).
For Question (1), we again find that both TD3 and SAC exhibit value and policy churn in all environments, and CHAIN-VCR and CHAIN-PCR effectively reduce them respectively in Figure 11 and 12. For Question (2), Figure 6 shows the evaluation results regarding episode return. We can see that CHAIN-PCR often improves the learning performance, especially for Ant-v4; in contrast, CHAIN-VCR improves slightly. We hypothesize that this is because the policy interacts with the environment directly, and the target critic-network also helps to reduce the value churn due to its delayed synchronization with the online critic.
Due to the limitation of space, we refer the readers to Appendix D.3 for more results on churn reduction, the influence of different choices of $\lambda_{Q},\lambda_{\pi}$, the results of combining VCR and PCR, as well as the effect of auto-adjustment of the regularization coefficient for TD3 and SAC.
@@
Offline AC methods In Offline RL, a policy is trained over a fixed dataset. We investigate if reducing churn can also improve the convergence of Offline RL. We use IQL (Kostrikov et al., 2022) with D4RL Antmaze dataset (Fu et al., 2020) and AWAC (Nair et al., 2020) with Adroit for our demonstration due to their popularity and good performance in corresponding tasks. Concretely, we use the public implementation and benchmark scores for IQL and AWAC from CORL [^2] https://github.com/tinkoff-ai/CORL. To apply CHAIN to IQL and AWAC, we implement the regularization for value churn reduction (VCR, Eq. 7) and policy churn reduction (PCR, Eq. 8) separately by adding a couple of lines of code without any other modification. We use $\lambda_{\pi}=1e3$ for both CHAIN IQL (PCR) and CHAIN AWAC (PCR); $\lambda_{Q}=0.01$ for CHAIN IQL (VCR) and $0.1$ for CHAIN AWAC (VCR) across different tasks. The results are summarized in Table 1 and 2.
@@
We observe that both CHAIN PCR and CHAIN VCR improve the scores for IQL and AWAC in most Antmaze and Adroit tasks. We hypothesize that CHAIN suppresses churn in the training of value and policy networks, thus reducing the bias caused by churn in parameter updates. One thing here that differs from TD3 and SAC considered in the online setting is that the policy network of IQL has no impact on the training of the value networks since the value networks (i.e., $Q$ and $V$) are trained purely based on in-sample data without accessing $a^{\prime}=\pi_{\phi}(s^{\prime})$. Thus, although IQL does not exhibit a chain effect explicitly, the policy and value networks of IQL still have churns, which are reduced by CHAIN in this case. We provide a further empirical study in Appendix D.4.
@@ 6.4 Scaling DRL Agents with CHAIN
It is widely known that scaling DRL agents up is challenging. Naively scaling DRL agents by widening or deepening the conventional MLP networks straightforwardly often fails and could even lead to collapse. Here, we investigate the relationship between churn and scale for DRL agents, as well as the effect of CHAIN on boosting scaling performance, to answer Question (3). We take PPO and MuJoCo tasks as the exemplary setting and scale up both the policy and value networks by a scale-up ratio within $\{2,4,8,16\}$ via widening or deepening. Note that the default network architecture (i.e., when the scale-up ratio equals one) for both the policy and value networks is a two-layer MLP with $256$ neurons for each layer, followed by an output layer.
As expected, we observed that the performance of PPO degraded severely as the scale-up ratio increased, as shown by the solid gray lines in Figure 7. Inspired by the prior study (Obando-Ceron et al., 2024), we found using a decreased learning rate as lr / sqrt(scale-up ratio) alleviates the degradation of PPO scaling to some degree (shown by the solid red lines). We then use the learning rate setting below by default. From the perspective of churn, we also observed that scaling PPO escalates the scale of the policy churn in PPO. More results can be found in Appendix D.5. Therefore, we then evaluate the effect of CHAIN in this scaling setting. The results are shown in Figure 7 with dashed lines. By comparing the lines in the same color, we found that CHAIN improves the learning performance of PPO across almost all scale-up ratios and the two learning rate settings.
Figure 7: The results regarding episode return for scaling PPO via widening. CHAIN helps to scale almost across all the configurations. Similar results can be found for widening scaling in Figure 20.
In addition, we extend the training horizon from 2M to 10M for Ant, Humanoid, and Walker2d. The results are reported in Table 3. For both widening or deepening cases, CHAIN helps to scale up PPO and achieves clear improvement in terms of episode return. Comparatively, scaling by widening slightly outperforms deepening, which echoes the observation in  (Ota et al., 2020) to some extent.
Table 3: Scaling PPO with CHAIN. Means and standard errors of final episode return over six seeds.
@@
Our results indicate that uncontrolled churn could be a possible reason for the scaling issue of DRL agents, and CHAIN improves scaling by reducing churn effectively. Though appealing, CHAIN does not fully address the scaling issue per se, and achieves sub-linear scaling on DRL agents.
7 Conclusion
In this paper, we conduct a formal study of churn in a general view and present the chain effect of value and policy churn. The chain effect indicates a compounding cycle, which biases parameter updates throughout learning. We propose an easy-to-implement method for value and policy churn reduction. Our experimental results demonstrate the effectiveness of our method in reducing churn and improving learning performance over a range of DRL environments and algorithms.
Acknowledgements
We want to acknowledge funding support from NSERC, FQRNT and CIFAR and compute support from Digital Research Alliance of Canada, Mila IDT, and NVidia.
Appendix A Limitations
Our work is limited in several directions below and we expect further studies on these points in the future.
  • First, the theoretical analysis of the chain effect under concrete assumptions remains to be explored. Other perspectives that may influence churn, such as network structure, representation learning, and experience replay are not considered in this work, which are worthwhile to study in the future.
  • Moreover, although we provided a simple method to adjust the regularization coefficients dynamically throughout learning by keeping a consistent relative loss scale, it is not sufficient for us to use the same hyperparameter for different domains. We believe that using normalization techniques to unify the scales in different domains is necessary to address this point, similar to the work in [Hafner et al., 2023].
  • Besides, for algorithms that involve the learning of both policy and value (e.g., deep AC methods), the implementation of CHAIN in different problems faces the question of choosing the best option among PCR, VCR and DCR. For this, we expect to develop a better method to integrate the effect of PCR and VCR in the future.
  • Another remaining problem is the lack of an in-depth understanding of the positive and negative effects of churn on the generalization of DRL agents, which could drive new methods that better leverage the potential of churn.
Appendix B Additional Formal Analysis
B.1 The NTK Expressions of Two Types of Deviation
@@
The NTK expression of the two types of deviation above indicates that action gradient deviation and policy value deviation are mainly influenced by the second-order partial derivatives of $Q_{\theta}$ and the deterministic policy gradient (DPG) of $\pi_{\phi}$ at the state $\bar{s}$ considered, in addition to the parameter update. An interesting observation here is, policy value deviation of policy churn has an implicit optimization effect (while altered by $\Delta_{\phi}$) for the reference states (i.e., $\bar{s}$) although the parameter update $\phi\rightarrow\phi^{\prime}$ is performed for the training states (i.e., $s$). We observe that $\mathcal{D}^{\pi}_{Q}$ is positive in overall in our empirical investigation later (Appendix D). However, careful considerations are needed in the future because the implicit optimization could be delusional as churn also occurs in $Q_{\theta}$ through the learning process.
B.2 Complete Discussions and Derivations for Section 4.3
@@
Appendix C Experimental Details
C.1 Compute Resources and Time Cost
@@
C.2 Empirical Metrics for the Investigation of the Chain Effect
@@
We use $N=50$ and $N=20$ for MuJoCo and MinAtar environments and compute the metrics at an interval of 1k parameter updates. For each type of metric in Table 4 and update number $i\in\{1,\dots,N\}$, we compute the mean of the quantities throughout learning. We refer the readers to Figure 9, 11 and 12 for the results in Appendix D.
C.3 Code Implementation
@@ C.4 Other Discussions
Churn Reduction Hyperparameters (refer to the study in Fig. 6, 11, 12)
@@
Appendix D Complete Results
D.1 More Empirical Analysis on the Value Churn in DoubleDQN
@@
D.2 More Results of CHAIN PPO
@@
D.3 More Results of CHAIN for Deep AC Methods
@@
Moreover, we observe a positive policy value deviation of action churn ($\hat{\mathcal{D}}^{\pi}_{Q}$), which matches our discussion on implicit optimization effect of the policy churn in Appendix B; in contrast, the value churn ($\hat{\mathcal{C}}_{Q_{sa}}$) fluctuates above and below 0.
@@
Figure 16 shows the evaluation of the automatic adjustment method of $\lambda_{\pi}$ for CHAIN TD3 in four MuJoCo tasks. We can observe that CHAIN (Auto) achieves comparable performance with manually selected coefficients and improves TD3, similarly to the conclusions we found in Figure 4 and 5.
@@
D.4 More Results of CHAIN for IQL with Sequential Training
@@
We report the final scores of IQL (sequential) with means and standard errors over 12 seeds in Table 8. The results show that IQL (sequential) performs worse than IQL in 5 of 6 tasks. The difference between IQL (sequential) and IQL can be fully attributed to the difference in the training dynamics, mainly on the policy network. This also means exposing the actor network to the training process of the $Q$ network is helpful compared to training the actor based on a sufficiently trained and frozen $Q$ network. To provide some possible explanation, we guess that this is because the final $Q$ network is a product of accumulated value churns, and the difference in the training dynamics of the policy network of IQL results in a further difference in the policy output on out-of-sample states. However, it is somewhat tricky to explain the difference between IQL (sequential) and IQL. The dynamics is beyond the scope of the chain effect of churn mechanism studied in our work.
@@
D.5 More Results for Scaling PPO with CHAIN